OpenStudy (mikezack123):

Pamela's age is three times Jiri's age. The sum of their ages is 68. What is Jiri's age?

Vocaloid (vocaloid):

let p = Pamela's age let j = Jiri's age j+p = 68 p = 3j solve for j

OpenStudy (mikezack123):

j=p/3

OpenStudy (mikezack123):

right?

OpenStudy (jacob902):

p = 3j p + j = 64 4j = 64 j = 16

OpenStudy (jacob902):

Pamela's age = 3 x Jiri's age = x 3 x + x = 64 4 x = 64 x = 64/4 =16 So , Jiri's age = 16 years old :-

OpenStudy (mikezack123):

thank you both! :) :)

OpenStudy (mikezack123):

i wish i could give you both medals :/

OpenStudy (mikezack123):

lol 16 is the wrong answer

OpenStudy (mikezack123):

their combined age is 68 not 64 lol @jacob902

OpenStudy (mikezack123):

@Vocaloid could u please help me? :)

Vocaloid (vocaloid):

start back with our two equations j+p = 68 p = 3j since p = 3j, we can replace "p" with "3j" whenever we see it j + p = j + 3j = 4j = 48 solve for j

OpenStudy (mikezack123):

okay so 4j = 48 ?

OpenStudy (mikezack123):

j= 12

Vocaloid (vocaloid):

right

OpenStudy (mikezack123):

so 12 would be jiri's age?

OpenStudy (mikezack123):

or?

Vocaloid (vocaloid):

please look at the first thing I posted...

OpenStudy (mikezack123):

did i do something wrong?

OpenStudy (mikezack123):

i answered 12, it said it was wrong

OpenStudy (mikezack123):

:(

Vocaloid (vocaloid):

whoops, 4j = 68

OpenStudy (mikezack123):

j=17 ?

Vocaloid (vocaloid):

right, sorry about that > >

OpenStudy (mikezack123):

Thank you! :)